OpenStudy (anonymous):

Choose the value for x that makes the open sentence true. 5 + 4 ● x = 32 + x2 (Points : 1) 5 4 3 2

OpenStudy (anonymous):

is the dot for multiplication , and is that last part 2*x or x^2?

OpenStudy (anonymous):

yes the dot is for mutiplication and the last part is x^2

OpenStudy (anonymous):

acually the last part is x to the second power

OpenStudy (anonymous):

sorry, but for 5 + 4x = 32 + x^2, none of those answers fit

OpenStudy (anonymous):

what

OpenStudy (anonymous):

what answer did u get

OpenStudy (anonymous):

move the numbers to one side x^2 - 4x + 27(32-5) = 0 quadratic equation formula ax^2 + bx + c = 0 x = (-b (+/-) sqrt(b^2 - 4ac))/2a (plus or minus) a = 1, b = -4, c = 27 x = (4 + sqrt(16 - 108)) / 2 x = (4 - sqrt(16 - 108)) / 2 16 - 108 = -92 square root of a nefative number is "complex", signified by an "i" in the answer x = ( 4 + sqrt(92)i ) / 2 x = ( 4 - sqrt(92)i ) / 2 x = ( 4 - 9.59i ) / 2 x = ( 4 + 9.59i ) / 2 x = ( -5.59i ) / 2 x = ( 13.59i ) / 2 x = -2.79i or 6.795 (you always have 2 answers for quadratic equations)